# On the non-Markovian quantum control dynamics

### v3

@@ IV-B Feedback control with atomic operators
where $11\leq i,j\leq N-$ 1. In the following, we take the two-level atom case as an example.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\bar{\alpha}=-ig_{1}e^{-i\delta t}\bar{s}-\frac{F_{a}(t)\kappa}{2}\bar{\alpha}+\mathcal{E},$ (58a)
$\frac{\mathrm{d}}{\mathrm{d}t}\bar{s}=ig_{1}e^{i\delta t}\bar{\alpha}\bar{w}-F_{1}(t)\kappa_{1}\bar{s}+ig_{f}\left(\bar{\alpha}+\bar{\alpha}^{*}\right)\bar{w}$
$~{}~{}~{}~{}~{}~{}~{}~{}-g_{f}^{2}\left(\bar{s}-\bar{s}^{*}\right)+ig_{f}\bar{w}\xi(t),$ (58b)
$\frac{\mathrm{d}}{\mathrm{d}t}\bar{w}=-2ig_{1}e^{i\delta t}\bar{s}^{*}\bar{\alpha}+2ig_{1}e^{-i\delta t}\bar{s}\bar{\alpha}^{*}$
$-\left(F_{1}(t)+F_{1}^{*}(t)\right)\kappa_{1}\left(\bar{w}+1\right)-2ig_{f}\left(\bar{s}^{*}-\bar{s}\right)\left(\bar{\alpha}+\bar{\alpha}^{*}\right)$
$-2g_{f}^{2}\bar{w}-2ig_{f}\left(\bar{s}^{*}-\bar{s}\right)\xi(t),$ (58c)
@@ V Quantum control for the multiple Jaynes-Cummings model
$H_{M}$ $=\sum_{m=1}^{M}g_{m}\left(e^{-i\delta_{m}t}\sigma_{-}^{(m)}a^{{\dagger}}_{m}+e^{i\delta_{m}t}\sigma_{+}^{(m)}a_{m}\right)$ (60)
$~{}~{}~{}~{}+\sum_{m}J_{c}\left(a^{{\dagger}}_{m}a_{m+1}+a_{m}a^{{\dagger}}_{m+1}\right),$
@@
$\mathrm{d}\rho=\left\{-i\left[H_{M},\rho\right]+\sum_{m}\left[\left(\mathcal{F}_{m}^{*}+\mathcal{F}_{m}\right)\sigma_{-}^{(m)}\rho\sigma_{+}^{(m)}\right.\right.$ (61)
$\left.\left.-\mathcal{F}_{m}\sigma_{+}^{(m)}\sigma_{-}^{(m)}\rho-\mathcal{F}_{m}^{*}\rho\sigma_{+}^{(m)}\sigma_{-}^{(m)}+\kappa_{m}\mathcal{L}_{a_{m}}\left[\rho\right]\right]\right\}\mathrm{d}t,$
@@ V-A Dynamics without drives
$\dot{\mathbf{x}}_{m}=$ $\begin{bmatrix}-\mathcal{F}_{m}(t)\kappa^{(m)}&-ig_{m}e^{i\delta_{m}t}\\ -ig_{m}e^{-i\delta_{m}t}&-\kappa_{m}\end{bmatrix}\mathbf{x}_{m}$ (62)
$-\begin{bmatrix}0\\ iJ_{c}\end{bmatrix}\langle a_{m+1}\rangle-\begin{bmatrix}0\\ iJ_{c}\end{bmatrix}\langle a_{m-1}\rangle.$
@@
where $\mathbf{\mathcal{A}}^{F}_{m}(t)=\begin{bmatrix}\mathcal{F}^{R}_{m}(t)&-\mathcal{F}^{I}_{m}(t)\\ \mathcal{F}^{I}_{m}(t)&\mathcal{F}^{R}_{m}(t)\end{bmatrix}$, $\mathbf{\mathcal{A}}^{G}_{m}(t)=\begin{bmatrix}g_{m}\sin\left(\delta_{m}t\right)&g_{m}\cos\left(\delta_{m}t\right)\\ -g_{m}\cos\left(\delta_{m}t\right)&g_{m}\sin\left(\delta_{m}t\right)\end{bmatrix}$, $\mathbf{\mathcal{A}}^{g}_{m}(t)=\begin{bmatrix}-g_{m}\sin\left(\delta_{m}t\right)&g_{m}\cos\left(\delta_{m}t\right)\\ -g_{m}\cos\left(\delta_{m}t\right)&-g_{m}\sin\left(\delta_{m}t\right)\end{bmatrix}$, $\mathbf{\mathcal{A}}_{\kappa}=\begin{bmatrix}-\kappa&0\\ 0&-\kappa\end{bmatrix}$ and $\mathbf{\mathcal{A}}_{J}=\begin{bmatrix}0&J_{c}\\ -J_{c}&0\end{bmatrix}$.
@@
When the interaction between the atom and the environment is Markovian or converges from non-Markovian to Markovian interactions, $\mathcal{F}_{m}$ equal constants and $\mathbf{\mathcal{A}}_{M}$ in Eq. (63) is periodic as $\mathbf{\mathcal{A}}_{M}\left(t+2\pi/\delta_{m}\right)=\mathbf{\mathcal{A}}_{M}\left(t\right)$ when $\delta_{1}=\cdots=\delta_{M}\neq 0$. Besides, when $\delta_{1}=\cdots=\delta_{M}=0$, $\mathbf{\mathcal{A}}_{M}\left(t\right)$ reduces to be time invariant.
Obviously, the stability of the high-dimensional equation (63), which is linear time-varying, can be similarly clarified by Corollary 2. For example, when there are two coupled cavities, the dynamics based on Markovian interaction between the quantum system and the environment has been analyzed in Ref. [70]. In this case, we denote $\mathbf{\mathcal{X}}_{\mathbf{a}}=\left[\mathbf{a}^{R}_{1},\mathbf{a}^{I}_{1},\mathbf{a}^{R}_{2},\mathbf{a}^{I}_{2}\right]^{\rm T}$, as a subspace of $\mathbf{\mathcal{X}}$. Obviously, when $g_{m}=0$ for $m=1,2$,
@@
which is linear and time-invariant. $\mathbf{\mathcal{X}}_{\mathbf{s}}$ is exponentially stable because in Eq. (68) the real part of roots determined by $(\textit{s}+\kappa)^{2}+J_{c}^{2}=0$ are negative, where s is for the Laplace transformation. Further by Corollary 1, $\lim_{t\rightarrow\infty}\mathbf{\mathcal{X}}_{\mathbf{a}}(t)=0$. This can be further generalized to the circumstance that $M>2$.
@@ V-B Dynamics with drives
The drive field can be applied via the feedforward or feedback channel. Take the measurement feedback as an example, as a generalization of Section IV-A, the measurement information can be collected from an arbitrary cavity of the coupled cavity array, and the measurement information from the $m$-th cavity reads
$I_{c}^{(m)}(t)$ $=\sqrt{2}\left\langle x_{m}\right\rangle+\frac{1}{\sqrt{\eta}}\xi_{m}(t),$ (69)
@@
$\left\langle\dot{\sigma}_{-}^{(m)}\right\rangle=-ig_{m}e^{i\delta_{m}t}\left\langle a_{m}\right\rangle-\mathcal{F}_{m}(t)\kappa^{(1)}\left\langle\sigma_{-}^{(m)}\right\rangle,$ (70a)
$\left\langle\dot{a}_{m}\right\rangle=-i\Delta\left\langle a_{m}\right\rangle-ig_{m}e^{-i\delta_{m}t}\left\langle\sigma_{-}^{(m)}\right\rangle-\kappa_{m}\left\langle a_{m}\right\rangle$
$-ig_{f}\left(\beta_{x}^{(m)}+i\beta_{p}^{(m)}\right)\left\langle x_{m}\right\rangle-iJ_{c}\left(\left\langle a_{m-1}\right\rangle+\left\langle a_{m+1}\right\rangle\right).$ (70b)
@@
When $g_{m}=0$, $\left\langle\sigma_{-}^{(m)}\right\rangle$ can converge to zero, while $\left\langle a_{m}\right\rangle$ can be unstable if the feedback parameter $g_{f}$ is large enough compared with $\left|\mathcal{F}_{m}(t)\right|$ and $\kappa$.
@@
$\mathbf{\mathcal{A}}_{u}=$ $\begin{bmatrix}\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa&\Delta&\cdots&0\\ \sqrt{2}g_{f}\beta_{x}^{(1)}-\Delta&-\kappa&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\Delta\\ 0&0&\cdots&-\kappa\end{bmatrix}$ (72)
$+\begin{bmatrix}0_{2\times 2}&\mathbf{\mathcal{A}}_{J}&0_{2\times 2}&\cdots&0_{2\times 2}&0_{2\times 2}\\ \mathbf{\mathcal{A}}_{J}&0_{2\times 2}&\mathbf{\mathcal{A}}_{J}&\cdots&0_{2\times 2}&0_{2\times 2}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0_{2\times 2}&0_{2\times 2}&0_{2\times 2}&\cdots&0_{2\times 2}&\mathbf{\mathcal{A}}_{J}\\ 0_{2\times 2}&0_{2\times 2}&0_{2\times 2}&\cdots&\mathbf{\mathcal{A}}_{J}&0_{2\times 2}\end{bmatrix},$
@@
For a special case that $g_{m}=0$, $P_{2}(t)=P_{3}(t)=0$, the dynamics of $X_{u}$ and $X_{s}$ are decoupled in Eq. (71). Obviously, $X_{s}$ will converge to zero if only the interaction between the atom and the environment becomes Markovian. However, $X_{u}$ can be unstable because of the feedback.
@@
$\frac{1}{2}\lambda_{\max}\left(\begin{bmatrix}\mathbf{\mathcal{A}}_{u}+\mathbf{\mathcal{A}}_{u}^{\rm T}&\mathcal{Q}_{2}(t)+\mathcal{Q}_{3}^{\rm T}(t)\\ \mathcal{Q}_{3}(t)+\mathcal{Q}_{2}^{\rm T}(t)&\mathbf{\mathcal{A}}_{s}+\mathbf{\mathcal{A}}_{s}^{\rm T}\end{bmatrix}\right)$ (75)
$=$ $\frac{1}{2}\lambda_{\max}\left(\begin{bmatrix}\bar{\mathbf{a}}_{1}&\cdots&0&0&\cdots&0\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ 0&\cdots&\bar{\mathbf{a}}_{M}&0&\cdots&0\\ 0&\cdots&0&\bar{\mathbf{b}}_{1}&\cdots&0\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ 0&\cdots&0&0&\cdots&\bar{\mathbf{b}}_{M}\end{bmatrix}\right),$
where $\bar{\mathbf{a}}_{m}=\begin{bmatrix}2\sqrt{2}g_{f}\beta_{p}^{(m)}-2\kappa_{m}&\sqrt{2}g_{f}\beta_{x}^{(m)}\\ \sqrt{2}g_{f}\beta_{x}^{(m)}&-2\kappa_{m}\end{bmatrix}$, $\bar{\mathbf{b}}_{m}=\begin{bmatrix}-2\bar{R}_{f}^{(m)}&0\\ 0&-2\bar{R}_{f}^{(m)}\end{bmatrix}$. When $g_{f}$ and $g_{x}$ are small, the inequality in (74) can be easier to be satisfied, then both $X_{u}$ and $X_{s}$ can converge to zero. An extreme case is that $\beta_{x}^{(m)}=0$ and $\kappa_{m}>\sqrt{2}g_{f}\beta_{p}^{(m)}$, then the matrix in Eq. (LABEL:con:MPMtr) is a diagonal matrix with all the diagonal elements being negative.
@@
As compared in Fig. 6, there are two coupled cavities, $J_{c}=0.1$GHz, $\delta=0.2$GHz, $\kappa=1$GHz, $g_{1}=0.2$GHz, $g_{2}=0.4$GHz, $\Delta=5$GHz, the two atoms in the two cavities are coupled identically with the non-Markovian environment with $\omega_{a}=43.98$GHz, and $\Omega=38.98$GHz, $\chi=1$, $\kappa^{(1)}=\kappa^{(2)}=0.04$GHz $\gamma=2$GHz for the non-Markovian environment [53]. In the measurement feedback design, $\beta_{x}^{(1)}=\beta_{x}^{(2)}=0$GHz, $\beta_{p}^{(1)}=\beta_{p}^{(2)}=0.2$GHz. On one hand, when $g_{f}=1$, $\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa=-0.7172{\rm GHz}<0{\rm GHz}$. On the other hand, when $g_{f}=7$, $\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa=0.9799{\rm GHz}>0{\rm GHz}$. The two group of simulations agree with Remark 7, showing that the stable and unstable subspaces can be modulated by tuning the feedback parameters.
The above analysis on multiple Jaynes-Cummings model realized by cascade cavity networks can be a potential approach for realizing non-Markovian interactions, according to the introduction on the frequency spectrum properties in [71]. Besides, the coupled cavity array can be further used to realize waveguide quantum electrodynamics [72], constructing various complex non-Markovian quantum networks [10, 11, 13]. The above generalization can introduce more challenges on quantum control due to the fact that the non-Markovian environment can destruct the purity of quantum states [73], which can further influence the control properties such as the creation of entangled quantum states [20] and photons [13].

### v4

@@ IV-B Feedback control with atomic operators
where $1\leq i,j\leq N-1$. In the following, we take the two-level atom case as an example.
@@
$\dot{\bar{\alpha}}=-ig_{1}e^{-i\delta t}\bar{s}-F_{a}(t)\kappa\bar{\alpha}/2+\mathcal{E},$ (56a)
$\dot{\bar{s}}=ig_{1}e^{i\delta t}\bar{\alpha}\bar{w}-F_{1}(t)\kappa_{1}\bar{s}+ig_{f}\left(\bar{\alpha}+\bar{\alpha}^{*}\right)\bar{w}-g_{f}^{2}\left(\bar{s}-\bar{s}^{*}\right)+ig_{f}\bar{w}\xi(t),$ (56b)
$\dot{\bar{w}}=-2ig_{1}e^{i\delta t}\bar{s}^{*}\bar{\alpha}+2ig_{1}e^{-i\delta t}\bar{s}\bar{\alpha}^{*}-\left(F_{1}(t)+F_{1}^{*}(t)\right)\kappa_{1}\left(\bar{w}+1\right)-2ig_{f}\left(\bar{s}^{*}-\bar{s}\right)\left(\bar{\alpha}+\bar{\alpha}^{*}\right)-2g_{f}^{2}\bar{w}-2ig_{f}\left(\bar{s}^{*}-\bar{s}\right)\xi(t),$ (56c)
@@ V Quantum control for the multiple Jaynes-Cummings model
$H_{M}$ $=\sum_{m=1}^{M}g_{m}\left(e^{-i\delta_{m}t}\sigma_{-}^{(m)}a^{{\dagger}}_{m}+e^{i\delta_{m}t}\sigma_{+}^{(m)}a_{m}\right)+\sum_{m}J_{c}\left(a^{{\dagger}}_{m}a_{m+1}+a_{m}a^{{\dagger}}_{m+1}\right),$ (58)
@@
$\dot{\rho}=-i\left[H_{M},\rho\right]+\sum_{m}\left[\kappa^{(m)}\left(\mathcal{F}_{m}^{*}+\mathcal{F}_{m}\sigma_{-}^{(m)}\rho\sigma_{+}^{(m)}-\mathcal{F}_{m}\sigma_{+}^{(m)}\sigma_{-}^{(m)}\rho-\mathcal{F}_{m}^{*}\rho\sigma_{+}^{(m)}\sigma_{-}^{(m)}\right)+\kappa_{m}\mathcal{L}_{a_{m}}\left[\rho\right]\right],$ (59)
@@ V-A Dynamics without drives
$\dot{\mathbf{x}}_{m}=$ $\begin{bmatrix}-\mathcal{F}_{m}(t)\kappa^{(m)}&-ig_{m}e^{i\delta_{m}t}\\ -ig_{m}e^{-i\delta_{m}t}&-\kappa_{m}\end{bmatrix}\mathbf{x}_{m}-\begin{bmatrix}0\\ iJ_{c}\end{bmatrix}\langle a_{m+1}\rangle-\begin{bmatrix}0\\ iJ_{c}\end{bmatrix}\langle a_{m-1}\rangle.$ (60)
@@
with $\mathbf{\mathcal{A}}^{F}_{m}(t)=\begin{bmatrix}\mathcal{F}^{R}_{m}(t)&-\mathcal{F}^{I}_{m}(t)\\ \mathcal{F}^{I}_{m}(t)&\mathcal{F}^{R}_{m}(t)\end{bmatrix}$, $\mathbf{\mathcal{A}}^{G}_{m}(t)=\begin{bmatrix}g_{m}\sin\left(\delta_{m}t\right)&g_{m}\cos\left(\delta_{m}t\right)\\ -g_{m}\cos\left(\delta_{m}t\right)&g_{m}\sin\left(\delta_{m}t\right)\end{bmatrix}$, $\mathbf{\mathcal{A}}^{g}_{m}(t)=\begin{bmatrix}-g_{m}\sin\left(\delta_{m}t\right)&g_{m}\cos\left(\delta_{m}t\right)\\ -g_{m}\cos\left(\delta_{m}t\right)&-g_{m}\sin\left(\delta_{m}t\right)\end{bmatrix}$, $\mathbf{\mathcal{A}}_{\kappa}=\begin{bmatrix}-\kappa&0\\ 0&-\kappa\end{bmatrix}$ and $\mathbf{\mathcal{A}}_{J}=\begin{bmatrix}0&J_{c}\\ -J_{c}&0\end{bmatrix}$.
@@
When the interaction between the atom and the environment is Markovian or converges from non-Markovian to Markovian interactions, $\mathcal{F}_{m}$ equals a constant and $\mathbf{\mathcal{A}}_{M}$ in Eq. (61) is periodic as $\mathbf{\mathcal{A}}_{M}\left(t+2\pi/\delta_{m}\right)=\mathbf{\mathcal{A}}_{M}\left(t\right)$ when $\delta_{1}=\cdots=\delta_{M}\neq 0$. Besides, when $\delta_{1}=\cdots=\delta_{M}=0$, $\mathbf{\mathcal{A}}_{M}$ reduces to be time invariant.
Obviously, the stability of the high-dimensional equation (61), which is linear time-varying, can be similarly clarified by Corollary 2. For example, when there are two coupled cavities, the dynamics based on Markovian interactions between the quantum system and the environment has been analyzed in Refs. [60, 74]. In this case, we denote $\mathbf{\mathcal{X}}_{\mathbf{a}}=\left[\mathbf{a}^{R}_{1},\mathbf{a}^{I}_{1},\mathbf{a}^{R}_{2},\mathbf{a}^{I}_{2}\right]^{\rm T}$, as a subspace of $\mathbf{\mathcal{X}}$. Obviously, when $g_{m}=0$ for $m=1,2$,
@@
which is linear and time-invariant. $\mathbf{\mathcal{X}}_{\mathbf{s}}$ is exponentially stable because in the Laplace transformation based on Eq. (66), the real part of roots determined by $(\textit{s}+\kappa)^{2}+J_{c}^{2}=0$ are negative. Further by Corollary 1, $\lim_{t\rightarrow\infty}\mathbf{\mathcal{X}}_{\mathbf{a}}(t)=0$. This can be further generalized to the circumstance that $M>2$.
@@ V-B Dynamics with drives
The drive field can be applied via the feedforward or feedback channel. Here we take the measurement feedback as an example. Generalized from Sec. IV-A, the measurement information can be collected from an arbitrary cavity of the coupled cavity array, and the measurement information from the $m$-th cavity reads
$I_{c}^{(m)}(t)$ $=\sqrt{2}\left\langle x_{m}\right\rangle+\frac{\xi_{m}(t)}{\sqrt{\eta}},$ (67)
@@
$\left\langle\dot{\sigma}_{-}^{(m)}\right\rangle=-ig_{m}e^{i\delta_{m}t}\left\langle a_{m}\right\rangle-\mathcal{F}_{m}(t)\kappa^{(m)}\left\langle\sigma_{-}^{(m)}\right\rangle,$ (68a)
$\left\langle\dot{a}_{m}\right\rangle=-i\Delta\left\langle a_{m}\right\rangle-ig_{m}e^{-i\delta_{m}t}\left\langle\sigma_{-}^{(m)}\right\rangle-\kappa_{m}\left\langle a_{m}\right\rangle-ig_{f}\left(\beta_{x}^{(m)}+i\beta_{p}^{(m)}\right)\left\langle x_{m}\right\rangle-iJ_{c}\left(\left\langle a_{m-1}\right\rangle+\left\langle a_{m+1}\right\rangle\right).$ (68b)
@@
Upon Assumption 2, when $g_{m}=0$, $\left\langle\sigma_{-}^{(m)}\right\rangle$ can converge to zero, while $\left\langle a_{m}\right\rangle$ can be unstable if the feedback parameter $g_{f}$ is large enough compared with $\left|\mathcal{F}_{m}(t)\right|$ and $\kappa$.
@@
$\mathbf{\mathcal{A}}_{u}=$ $\begin{bmatrix}\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa&\Delta&\cdots&0\\ \sqrt{2}g_{f}\beta_{x}^{(1)}-\Delta&-\kappa&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\Delta\\ 0&0&\cdots&-\kappa\end{bmatrix}+\begin{bmatrix}0_{2\times 2}&\mathbf{\mathcal{A}}_{J}&0_{2\times 2}&\cdots&0_{2\times 2}&0_{2\times 2}\\ \mathbf{\mathcal{A}}_{J}&0_{2\times 2}&\mathbf{\mathcal{A}}_{J}&\cdots&0_{2\times 2}&0_{2\times 2}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0_{2\times 2}&0_{2\times 2}&0_{2\times 2}&\cdots&0_{2\times 2}&\mathbf{\mathcal{A}}_{J}\\ 0_{2\times 2}&0_{2\times 2}&0_{2\times 2}&\cdots&\mathbf{\mathcal{A}}_{J}&0_{2\times 2}\end{bmatrix}$, (70)
@@
For a special case that $g_{m}=0$, $P_{2}(t)=P_{3}(t)=0$, the dynamics of $X_{u}$ and $X_{s}$ are decoupled in Eq. (69). Obviously, $X_{s}$ will converge to zero if only the interaction between the atom and the environment becomes Markovian. However, $X_{u}$ can be unstable because of feedback controls.
@@
[…]
where $\bar{\mathbf{a}}_{m}=\begin{bmatrix}2\sqrt{2}g_{f}\beta_{p}^{(m)}-2\kappa_{m}&\sqrt{2}g_{f}\beta_{x}^{(m)}\\ \sqrt{2}g_{f}\beta_{x}^{(m)}&-2\kappa_{m}\end{bmatrix}$, $\bar{\mathbf{b}}_{m}=\begin{bmatrix}-2\bar{R}_{f}^{(m)}&0\\ 0&-2\bar{R}_{f}^{(m)}\end{bmatrix}$. When $g_{f}$ and $\beta_{x}^{(m)}$ are small, the inequality in (72) can be easier to meet, then both $X_{u}$ and $X_{s}$ can converge to zero. An extreme case is that $\beta_{x}^{(m)}=0$ and $\kappa_{m}>\sqrt{2}g_{f}\beta_{p}^{(m)}$, then the matrix in Eq. (LABEL:con:MPMtr) is a diagonal matrix with all the diagonal elements being negative.
@@
As compared in Fig. 6, there are two coupled cavities, $J_{c}=0.1$GHz, $\delta=0.2$GHz, $\kappa=1$GHz, $g_{1}=0.2$GHz, $g_{2}=0.4$GHz, $\Delta=5$GHz, the two atoms in the two cavities are coupled identically with the non-Markovian environment with $\omega_{a}=43.98$GHz, and $\Omega=38.98$GHz, $\chi=1$, $\kappa^{(1)}=\kappa^{(2)}=0.04$GHz, $\gamma=2$GHz for the non-Markovian environment [55]. In the measurement feedback design, $\beta_{x}^{(1)}=\beta_{x}^{(2)}=0$GHz, $\beta_{p}^{(1)}=\beta_{p}^{(2)}=0.2$GHz. On one hand, when $g_{f}=1$, $\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa=-0.7172{\rm GHz}<0{\rm GHz}$. On the other hand, when $g_{f}=7$, $\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa=0.9799{\rm GHz}>0{\rm GHz}$. The two group of simulations agree with Remark 8, showing that the stable and unstable subspaces can be modulated by tuning the feedback parameters.
The above analysis on multiple Jaynes-Cummings model realized by cascade cavity networks can be a potential approach for realizing non-Markovian interactions, according to the introduction on the frequency spectrum properties in [75]. Besides, the coupled cavity array can be further used to realize waveguide quantum electrodynamics [76], constructing various complex non-Markovian quantum networks [10, 11, 12]. This generalization can introduce more challenges on quantum control due to the fact that the non-Markovian environment can destruct the purity of quantum states [77], which can further influence the control properties such as the creation of entangled quantum states [19] and photons [12].